OpenStudy (jtrippz):

if y=3 when x=9 find x when y=1 & if y=2 when x=0.5 find y when x=3

OpenStudy (jonathan34):

hi

OpenStudy (sophiesky):

\[\frac{ 9 }{ 3 }=\frac{ x }{ 1 }\] Now cross multiply and solve for x.

OpenStudy (sophiesky):

Do you know how to cross multiply?

OpenStudy (jtrippz):

Yes but I also need the answer for the second problem

OpenStudy (sophiesky):

Alright, well what's the answer to the first problem?

OpenStudy (jtrippz):

x=3 ??? lol

OpenStudy (sophiesky):

Correct! Here's the second problem. \[\frac{ .5 }{ 2 }=\frac{ 3 }{ y }\]

OpenStudy (jtrippz):

and do the same for the previous problem?

OpenStudy (sophiesky):

Yes!

OpenStudy (jtrippz):

y=1.2 ? I calculated and that's the answer I got

OpenStudy (sophiesky):

I think y is 12..

OpenStudy (jtrippz):

either one will work lol

OpenStudy (sophiesky):

\[\frac{ 6 }{ .5 }=12\]

OpenStudy (jtrippz):

Oh okay, thanks!

OpenStudy (sophiesky):

Of course! Need anymore help?

OpenStudy (jtrippz):

Can you also help me with y=2 when x=1 and y=-3 when x=1

OpenStudy (sophiesky):

So usually it's \[\frac{ x }{ y }\] and so how about you try this problem?